estimate the number of molecules in 1.00m^3 of air by assuming that it has a density of 1.21kg/m^3 and that is entirely composed of nitrogen molecules


1
Expert's answer
2021-02-07T19:10:12-0500

M(N2) = 28.014 g/mol

ρ(N2) = 1.21 kg/m3

V(N2) = 1 m3

"m = ρ \times V \\

m(N_2) = 1.21 \times 1 = 1.21 \;kg = 1210 \;g \\

n=\frac{m}{M} \\

n(N_2) = \frac{1210}{28.014} \\

= 43.19 \;mol \\

N_A = 6.022 \times 10^{23} \\

n = 43.19 \times 6.022 \times 10^{23} \\

= 2.6 \times 10^{25}"

Answer: "2.6 \times 10^{25}"